A company obtained two loans totaling $30,000. The interest rate for one loan was 16 percent, and for the other loan it was 12 percent. The total annual interest for the two loans was equal to the interest the company would have paid if the entire amount had been borrowed at 15 percent. What was the amount of each loan?
Answer by mananth(16949)

15% of 30,000= 4500
Investment I 16.00% per annum ---x
Investment II 12.00% per annum ---y

x + y= 30000 ------------------------1
16.00% x + 12.00% y= = $4,500.00
Multiply by 100
16 x + 12 y= = $450,000.00 --------2

Multiply (1) by -16
we get

-16 x -16 y= = -480000.00

Add this to (2)

-4 y= = -30,000.00

divide by -4

y = $7,500.00 investment at 12.00%
Balance $22,500.00 investment at 16.00%
CHECK
$22,500.00 --------- 16.00% ------- $3,600.00
$7,500.00 ------- 12.00% ------- $900.00
Total -------------- $4,500.00
m.ananth@hotmail.ca